Ask
your own question, for FREE!
Mathematics
12 Online
A card is randomly selected from a standard deck. What is the probability that a club or a red queen is selected? a 1/26 b 14/52 c 15/52 d 17/52
Still Need Help?
Join the QuestionCove community and study together with friends!
There are 13 clubs in the deck. Of the queens in the deck, 2 are red. Prob(Club or red queen) = 15/52
probability of a club, A, K, Q, J, 2, 3, 4,....10, would be 13/52 probability of a red queen, diamond or heart, would be 2/52 therefore 13/52 + 2/52 = 15/52 on "or" probability problems, one adds the two probabilities together
thanks! both of u!!
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
6 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal